### 07 Feb 2024 v1

Abstract
Diffusion models have enabled remarkably high-quality medical image generation, which can help mitigate the expenses of acquiring and annotating new images by supplementing small or imbalanced datasets, along with other applications. However, these are hampered by the challenge of enforcing global anatomical realism in generated images. To this end, we propose a diffusion model for anatomically-controlled medical image generation. Our model follows a multi-class anatomical segmentation mask at each sampling step and incorporates a random mask ablation training algorithm, to enable conditioning on a selected combination of anatomical constraints while allowing flexibility in other anatomical areas. This also improves the network’s learning of anatomical realism for the completely unconditional (unconstrained generation) case. Comparative evaluation on breast MRI and abdominal/neck-to-pelvis CT datasets demonstrates superior anatomical realism and input mask faithfulness over state-of-the-art models. We also offer an accessible codebase and release a dataset of generated paired breast MRIs. Our approach facilitates diverse applications, including pre-registered image generation, counterfactual scenarios, and others.
keywords:
Diffusion models, image generation, breast MRI, CT
Introduction
Denoising diffusion probabilistic models [Ho et al.(2020)Ho, Jain, and Abbeel] (DDPMs, or just “diffusion models”) have shown extensive promise in medical image analysis and beyond [Kazerouni et al.(2023)Kazerouni, Aghdam, Heidari, Azad, Fayyaz, Hacihaliloglu, and Merhof] due to their ability to generate high-quality, high-resolution images. However, standard generative models like DDPMs can still fail to create anatomically plausible tissue (Fig LABEL:fig:eg_imgs). Our proposed solution is to incorporate anatomical information as a prior for image generation via a segmentation mask for different types of tissue, organs, etc., providing the network with a more direct learning signal for anatomical realism. Image-space-based diffusion models are especially amenable for strict segmentation guidance because a conditioning mask can be used as-is for every small step of the denoising process, with no conversion to some abstract latent space as in latent diffusion models (or GANs) where complex spatial conditions may be lost. In this paper, we introduce an image-space diffusion model for segmentation-guided image generation: Segmentation-Guided Diffusion.
[…]
fig:eg_imgs Refer to caption
Figure 1: Standard diffusion models (right) can fail to create realistic tissue, even if the overall image appears high-quality. From left to right: (1) randomly sampled real breast MRI images, (2) synthetic images created by our segmentation-guided diffusion model, and (3) by the same model but without mask guidance; insets show areas with fibroglandular tissue.
Segmentation-guided generation would be even more flexible if only certain object classes could be constrained in an input mask, while others are free to be inferred/“filled in” by the model (also allowing the model to be applied to incomplete segmentations [Tian et al.(2023)Tian, Yu, and Fu]). This also opens up further applications such as the generation of anatomically paired/registered data, “counterfactual” analysis of existing annotated data, and others. To solve this, we propose a mask-ablated training strategy to provide the model with all possible combinations of missing classes in masks during training, teaching it to make such inferences when generating new images. Notably, we also show that this training strategy helps the model learn to generate more realistic tissue in the case of completely unconditional (empty mask) inference (Sec. 3.2), allowing for the “best of both worlds” of both anatomical accuracy and diverse and flexible unconditional generation.
We evaluate our model’s ability to generate images conditioned on a range of anatomical objects of interest for breast MRI and neck-to-pelvis CT datasets and compare it to state-of-the-art mask-conditional and unconditional generative models in terms of anatomical realism and faithfulness to input masks of generated images. Notably, our approach far outperforms the recent mask-conditional latent diffusion model ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala] in both metrics. We also publicly release our code at https://github.com/mazurowski-lab/segmentation-guided-diffusion with a focus on usability on any dataset of choice, along with a dataset of paired “pre-registered” generated breast MRIs.
Related Work
Diffusion Models for Medical Image Analysis
Image Generation from Spatial Conditions
Generating an image from a mask can be considered as an image-to-image translation or conditional generation task, where an input image is typically converted to some latent representation from which the image in the desired output space is decoded. Examples include GAN-based models [Yang et al.(2019)Yang, Dvornek, Zhang, Chapiro, Lin, and Duncan, Choi et al.(2018)Choi, Choi, Kim, Ha, Kim, and Choo, Wang et al.(2018)Wang, Liu, Zhu, Tao, Kautz, and Catanzaro], or diffusion-based models [Wolleb et al.(2022a)Wolleb, Bieder, Sandkühler, and Cattin] which partially noise the image, then guide the denoising process towards the target domain via various forms of image class guidance [Dhariwal and Nichol(2021), Ho and Salimans(2021)]. As opposed to our approach, these methods do not directly incorporate pixel-wise spatial guidance in the image generation process.
ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala] is a recent method for adding various fine-tuned spatial conditioning controls to Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], a large pre-trained text-to-image latent diffusion model. While ControlNet relies on first encoding an input mask to some latent space before feeding it to the diffusion model, our model always receives mask signals from image space directly. We find that this allows for more transparency and faithfulness with how the network can follow the mask conditions, and better anatomical realism (Sec. 3.1), and that ControlNet is generally challenging to adapt to medical images for various reasons (Appendix B).
1 Method
1.1 A Brief Review of Diffusion Models
Denoising diffusion probabilistic models [Ho et al.(2020)Ho, Jain, and Abbeel] (DDPMs, or diffusion models for short) are a type of generative latent variable model (LVM) that learns to sample from some data distribution $p(x_{0})$ ($x_{0}\in\mathbb{R}^{n}$) by defining a stochastic process that gradually converts the data to noise—the forward process $q(x_{t}|x_{t-1})$, constructed of conditional Gaussians—and learning to reverse this process via a learned denoising process $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ is the model parameters. Data is generated by iteratively sampling from $p_{\theta}(x_{t-1}|x_{t})$, beginning with a Gaussian noise sample $x_{T}\sim p(x_{T})$, for $t=T-1,\ldots,0$ until an image $x_{0}$ is recovered.
Any step of the forward process can be written explicitly as $x_{t}=\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(0,I_{n})$, and $\alpha_{t}:=1-\beta_{t}$ given the variance of the additive pre-scheduled noise $\beta_{t}$, and $\overline{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$. As is typical for LVMs [Kingma and Welling(2013)], DDPMs can be trained by maximizing the evidence lower bound (ELBO) with respect to $\theta$ given observed data $x_{0}$ as $\mathrm{ELBO}(\theta;x_{0})=\mathbb{E}_{x_{1:T}\sim q(x_{1:T}|x_{0})}\left[\log p(x_{T})+\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t}|x_{t-1})}\right]$, where $x_{1:T}:=x_{1},\ldots x_{T}$. The ELBO can be approximately optimized in a relatively simple form by training a network $\epsilon_{\theta}(x_{t},t)$ to predict the noise $\epsilon$ added to each datapoint $x_{0}$ for various time steps $t$, with the loss $L=\mathbb{E}_{x_{0},t,\epsilon}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}\right]=\mathbb{E}_{x_{0},t,\epsilon}\left[||\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,t)||^{2}\right]$, which has proven to be the typically superior DDPM loss in practice [Nichol and Dhariwal(2021)].
1.2 Adding Segmentation Guidance to Diffusion Models
Rather than sampling from the unconditional distribution $p(x_{0})$, our goal is to condition the generation of some image $x_{0}\in\mathbb{R}^{c\times h\times w}$ to follow some multi-class anatomical mask $m\in\{0,\ldots,C-1\}^{h\times w}$ (i.e., sample from $p(x_{0}|m)$), where $c$ is the number of channels of $x_{0}$ and the mask has $C$ possible values of $10,\ldots,C-$ 1 associated with each object class. While this does not alter the noising process $q(x_{t}|x_{t-1})$, the data log-likelihood becomes mask-conditioned as $\log p(x_{0}|m)$, thereby similarly modifying the reverse process $p_{\theta}(x_{t-1}|x_{t},m)$ and noise-predicting network $\epsilon_{\theta}$, changing the ELBO to
$\mathrm{ELBO}(\theta;(x_{0},m))=\mathbb{E}_{x_{1:T}\sim q(x_{1:T}|x_{0})}\left[\log p(x_{T})+\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}|x_{t},m)}{q(x_{t}|x_{t-1})}\right].$ (1)
We propose to implement this with a simple approach of concatenating the mask $m$ channel-wise to the network input at every denoising step, written as $\epsilon_{\theta}(x_{t},t|m)$. We modify the network to take an additional channel allocated for $m$ as input, i.e., $\epsilon_{\theta}:\mathbb{R}^{(c+1)\times h\times w}\rightarrow\mathbb{R}^{c\times h\times w}$, which can be any image-to-image model (see Sec. 3). This results in a loss of
$L_{m}=\mathbb{E}_{(x_{0},m),t,\epsilon}\left[||\epsilon-\epsilon_{\theta}(x_{t},t|m)||^{2}\right]=\mathbb{E}_{(x_{0},m),t,\epsilon}\left[||\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,t|m)||^{2}\right]$ (2)
for training our model, where each training image $x_{0}$ has some paired anatomical mask $m$. Sampling can be completed with the typical DDPM algorithm [Ho et al.(2020)Ho, Jain, and Abbeel] if stochastic results are desired given some input mask, although in practice we use the DDIM (denoising diffusion implicit model) algorithm [Song et al.(2021)Song, Meng, and Ermon], which provides 20×20$\times$20 × faster sampling than the standard DDPM algorithm with little reduction in visual quality.
This simple method generates images that are very faithful to input masks (Fig. LABEL:fig:samples), because the denoising process is conditioned on the mask at each of its many gradual steps, allowing the network to follow the masks because they provide helpful spatial information that is directly correlated with the optimal denoised model output that minimizes the loss.
1.3 Mask-Ablated Training and Sampling
Given that our model is mask-guided, the quality of these masks is important; a generated image may be misleading if the input mask is not fully annotated, known as the partial label problem [Tian et al.(2023)Tian, Yu, and Fu] in medical image analysis. This is because the model may assume that un-annotated objects should not be present in the output image whatsoever (associating the missing/zero pixel labels as background or some other object), when in reality we may desire for the model to simply “fill in”/infer the unprovided objects.
To alleviate this problem, we propose a mask-ablated training strategy (Algorithm 2), which provides examples of masks with various numbers and combinations of classes removed for the model to learn to generate images from during training. We set all $2^{C-1}$ of these possible combinations of classes being removed from a given mask in training (including the cases of no classes or all classes removed) to occur with equal probability so that the model can handle each equally, although any other balancing of these probabilities could be used.
[…]
Dodowhile
[…]
Figure 2: Segmentation-guided diffusion model training with mask ablation.
2 Datasets
Breast MRI: Our first dataset is a 100-patient subset of the Duke Breast Cancer MRI dataset [Saha et al.(2018)Saha, Harowicz, Grimm, Kim, Ghate, Walsh, and Mazurowski]. We use all 2D image slices from the fat-saturated gradient echo T1-weighted pre-contrast sequence, with a train/test split of 70/15 patients, resulting in $\sim$12000/2500 slice images per split. We also keep a held-out training set of 15 patients for additional experiments. All images have full segmentation annotations for (1) the breast, (2) blood vessels, and (3) fibroglandular/dense tissue provided at [Lew(2023)]. Notably, the fibroglandular tissue and blood vessels have very high variability in shape, size, and other morphological characteristics, posing a challenge for generative models to realistically capture. CT Organ: Our second dataset is CT-Organ [Rister et al.(2020)Rister, Yi, Shivakumar, Nobashi, and Rubin]. We use a 40-patient subset of neck-to-pelvis and abdominal CT scans, which include segmentation annotations for liver, bladder, lungs, kidney, and bone. This results in a train/test split of $\sim$11000/2100 2D slice images, given a patient-wise split of 24/8, as well as a held-out training set of 8 patients. All generative models are trained on the training sets, and the auxiliary segmentation network, introduced next, is trained on the held-out training sets.
3 Experiments
Training, Architecture and Implementational Details
All images are resized to $256\times 256$ and normalized to $[0,255]$. We use a UNet architecture [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] for the denoising model $\epsilon_{\theta}$, modified to take two channels (image and mask) as input; see Appendix A.1 for additional training and architecture details.
3.1 Comparison to Existing Image Generation Models
In this section, we will compare our segmentation-guided diffusion model (“Seg-Diff” for short) to various state-of-the-art image generation models: (1) the segmentation-conditional model ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala], and the unconditional models of (2) a diffusion model identical to ours but trained without mask conditioning, and (3) StyleGAN2-ADA [Karras et al.(2020)Karras, Aittala, Hellsten, Laine, Lehtinen, and Aila], a state-of-the-art GAN model. We abbreviate these as CtrlNet, Unc-Diff and GAN, respectively, and training details for each are in Appendix A.2. We show exemplar generated images from all models in Fig. LABEL:fig:samples; for conditional models, input masks were randomly sampled from the test set.
[…]
fig:samples Refer to caption Refer to caption
Figure 3: Randomly sampled breast MRI (top) and CT organ (bottom) images generated with different segmentation-conditional and unconditional models. Conditional model outputs are below the corresponding input mask. For breast MRI, the breast, blood vessel, and fibroglandular tissue segmentations are shown as white, red, and blue, respectively, while for CT, the liver, bladder, lungs, kidneys, and bone are in brown, orange, pink, red, and white, respectively.
Evaluating Generated Image Quality
We first attempted to use the common Fréchet Inception Distance (FID) [Heusel et al.(2017)Heusel, Ramsauer, Unterthiner, Nessler, and Hochreiter] as a metric for generated image quality/realism compared to real data. Since the original FID utilizes a CNN image encoder pre-trained on natural images, which may fail to capture the characteristics of medical images, we propose to instead use an encoder trained on the corresponding medical image dataset. We observe that samples generated from both the segmentation-guided and unconditional diffusion models achieve promising results, (e.g., $\mathrm{FID}\simeq 0.5$ for breast MRI), yet the CNN-based FID fails to capture the global feature of anatomical realism that can differ in images generated by these models (e.g., fibroglandular tissue as shown in Fig. LABEL:fig:eg_imgs).
Table 1: Quality of generated images. Real test set performance (Dice coeff.) of a segmentation network for the objects of interest, trained on real data vs. synthetic data generated by different models.
tab:qual
Training set Breast MRI CT Organ
Real 0.8376 0.9075
Synthetic:
 Our model 0.7934 0.8981
 ControlNet 0.7570 0.0000
Table 1: Quality of generated images. Real test set performance (Dice coeff.) of a segmentation network for the objects of interest, trained on real data vs. synthetic data generated by different models.
We propose to better measure anatomical realism by instead determining how well an auxiliary segmentation model for the objects of interest can be trained solely on these synthetic images to be able to generalize to real data, using the input masks as targets (training details in Appendix A.3). We compare the performance of the segmentation model trained on (a) the real held-out training set (Sec. 2) and (b) the set of images generated from all masks corresponding to these images. We split the real test set in half (by patient) into a validation set and a test set to use for these models. The results for this are in Table LABEL:tab:qual; we see that for both datasets, the segmentation network trained on our model’s synthetic data barely performs worse ($<0.04-0.01$ Dice) than the network trained on real data, implying that our generated images are both highly realistic and faithful to the input masks (especially considering certain objects’ segmentation difficulty), while outperforming ControlNet.
Table 2: Faithfulness of generated images to input masks. $m$ denotes generative model input masks, and $m^{pred}_{gen}$ and $m^{pred}_{real}$ denote the masks predicted for (a) the generated images or (b) the real images corresponding to the input masks, respectively, by an auxiliary segmentation model.
tab:faithful
Breast MRI CT Organ
Model $\mathrm{Dice}(m^{pred}_{gen},m)$ $\mathrm{Dice}(m^{pred}_{gen},m^{pred}_{real})$ $\mathrm{Dice}(m^{pred}_{gen},m)$ $\mathrm{Dice}(m^{pred}_{gen},m^{pred}_{real})$
Segmentation-guided
diffusion (ours)
0.9027 0.8593 0.8980 0.8797
ControlNet 0.3636 0.3604 0.1132 0.1126
Table 2: Faithfulness of generated images to input masks. $m$ denotes generative model input masks, and $m^{pred}_{gen}$ and $m^{pred}_{real}$ denote the masks predicted for (a) the generated images or (b) the real images corresponding to the input masks, respectively, by an auxiliary segmentation model.
Evaluating Faithfulness of Generated Images to Input Masks
To measure how well our model follows an input mask for image generation, we use the aforementioned segmentation network trained on the real training set, to predict segmentations $m^{pred}_{gen}$ on images that were generated from the masks $m$ in the test set, and measure their overlap with (a) $m$ and (b) the model’s predicted segmentations $m^{pred}_{real}$ for the input masks’ original corresponding real images. We see our model’s generated images have high overlap with both ($>0.85$ Dice coeff., Table LABEL:tab:faithful), indicating that our model closely followed the input masks when generating the images, also far outperforming ControlNet’s generated images.
Understanding ControlNet’s Performance
ControlNet performed much worse for each metric because it failed to generate images that follow input masks closely (if at all). We suspect that is because (1) their model’s ability to adapt to out-of-distribution (e.g., medical) images was untested in their paper, and (2) their training procedure for learning segmentation guidance requires joint image+text embeddings, the latter of which is not immediately applicable to our dataset. We discuss this in more detail in Appendix B.
[…]
fig:samples_ablated Refer to caption
Figure 4: Generating images from masks with classes removed, shown for breast MRI. Rows 1 and 3: input segmentations, Rows 2 and 4: corresponding output images.
3.2 Sampling from Ablated Masks
We will next explore our model’s ability to be conditioned on a range of combinations of present or missing anatomical classes in input masks, following our mask-ablated training algorithm (Sec. 1.3). In Fig. LABEL:fig:samples_ablated we demonstrate the effect on generated images of ablating certain classes from an input mask for breast MRI. For example, we see that constraining the blood vessels and fibroglandular tissue (BV+FGT) in breast MRI, yet keeping the breast free, results in images that have the latter two classes pre-registered, while the breast shape varies. Similarly, we see that constraining the breast shape while letting (BV+FGT) be free results in images with the same breast shape, yet different BV+FGT.
Notably, we also see that when the model is provided with an empty mask (unconditioned), it typically generates more anatomically realistic images than the unconditionally-trained model does (e.g., compare the rightmost image in Fig. LABEL:fig:samples_ablated to the rightmost image in Fig. LABEL:fig:eg_imgs), due to the ablated mask guidance that it learned from in training. This is important because the model has both the anatomical realism of a model trained with full mask guidance and the ability to generate new diverse samples without needing input masks.
Conclusion
Our segmentation-guided diffusion model enables more realistic medical image generation, as well as applications such as the synthesis of paired/pre-registered data. We also hope that our public codebase and generated dataset of paired breast MRIs provide helpful resources for the community. For future work, we are interested in incorporating image-level class guidance [Ho and Salimans(2021)] either for pathological or domain-related variables [Konz and Mazurowski(2023)], extending our model to segmentation-guided image translation, and extending to 3D imaging.
Appendix A Additional Training and Architectural Details
A.1 Our model (segmentation-guided diffusion)
The denoising model (UNet)’s encoder is constructed from six standard ResNet down-sampling blocks, with the fifth block also having spatial self-attention, with $(128,128,256,256,512,512)$ output channels, respectively. The decoder is simply the up-sampling reverse of the encoder. We use $T=1000$ timesteps for the denoising and noising processes, with a standard forward process variance schedule that linearly increases from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$ [Ho et al.(2020)Ho, Jain, and Abbeel]. For training, we use the AdamW optimizer [Loshchilov and Hutter(2018)] and a cosine learning rate scheduler [Loshchilov and Hutter(2016)] with an initial learning rate of $10^{-4}$, with $500$ linear warm-up steps. We train for 400 epochs with a batch size of 64, and we perform all training and evaluation on four 48 GB NVIDIA A6000 GPUs. We use the Diffusers library as a backbone [von Platen et al.()von Platen, Patil, Lozhkov, Cuenca, Lambert, Rasul, Davaadorj, and Wolf].
A.2 Comparison models
Unconditional DDPM.
The unconditional DDPM that we compare to has identical training and architecture to our segmentation-guided model, just without taking segmentations as input during training.
StyleGAN2-ADA.
We train StyleGAN2-ADA for $2$ million image iterations (approximately 200 epochs), a batch size of 64, and all other settings set to the auto-generated default of https://github.com/NVlabs/stylegan2-ada-pytorch (except with the network set for one-channel images.
ControlNet.
We adapted ControlNet to each of our medical image datasets as was instructed at their official tutorial for use with datasets that are out-of-distribution (e.g., medical images) from the Stable Diffusion (SD) model’s very large natural image training set (https://github.com/lllyasviel/ControlNet/blob/main/docs/train.md#sd_locked). This involved first finetuning the autoencoder for 200 epochs, then finetuning the Stable Diffusion (SD) model for 400 epochs using the respective breast MRI or CT organ training set images. We then finetuned the ControlNet with the images and their corresponding masks for segmentation guidance, for 200 epochs, with empty text prompts. The pretrained (pre-finetuning) models are from the SD v1.5 checkpoints available on Hugging Face at https://huggingface.co/runwayml/stable-diffusion-v1-5. For all training, we set the batch size to 128, the initial learning rate to $10^{-4}$, and adopted cosine annealing learning rate schedulers rate with 500 steps of warm-up.
We note that our model being image-space-based, instead of a latent diffusion model, also allows for faster and simpler training-from-scratch on a new dataset; for example, on breast MRI we found that our model took 4-5 days to train on one 48 GB A6000, while ControlNet took 8-9 days (including the autoencoder, SD/latent diffusion model, and the ControlNet itself).
A.3 Auxiliary segmentation model
We used the UNet segmentation model from MONAI (https://docs.monai.io/en/stable/networks.html) with single channel input and the number of output channels set to the number of target object classes + 1 (background). The sequence of intermediate UNet channels was set to (16, 32, 64, 128, 256). For training, we set the initial learning rate to $10^{-3}$ and adopted a cosine annealing scheduler. The input image resolution was $256\times 256$. We trained each model for 100 epochs with a batch size of 8 and selected the models with the lowest validation loss.
Appendix B Challenges of Adapting ControlNet for Medical Images
As shown in our main results, ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala] failed to achieve solid segmentation-guided performance when adapted to our datasets, which we believe to be due to several reasons.
First, as mentioned, we adapted ControlNet to each of our medical image datasets as was instructed at their official tutorial for use with datasets that are out-of-distribution (e.g., medical images) from the Stable Diffusion (SD) model’s very large natural image training set (https://github.com/lllyasviel/ControlNet/blob/main/docs/train.md#sd_locked). This involved (1) fine-tuning the latent space VAE, (2) fine-tuning the SD model and (3) training the ControlNet module to learn segmentation guidance, on the new given dataset. However, we note that this procedure of finetuning the SD model was completely untested in their paper, and so was unproven. ControlNet works by first copying the SD model weights to create the ControlNet module itself, and then fine-tuning that module for spatial conditioning. However, this was not evaluated (or proposed) for a custom dataset finetuned SD model, so the performance being unreliable is understandable.
Second, the ControlNet training algorithm requires using non-empty text prompts for half of the images seen in training, while in our case, no text exists for our data, such that we had to train with all text prompts empty; as the diffusion model acts completely on the joint image-text latent space, the effect of this difference in training on the latent space is unclear and possibly violates the assumptions needed for ControlNet to work.
We also note that we performed each of these trainings for many epochs (Appendix B), ruling out the possible issue of the “sudden convergence phenomenon” of proper segmentation conditioning that was observed in the ControlNet paper. Finally, we could foreseeable train some latent diffusion model from scratch on each of our datasets (without text conditioning), but this is out-of-scope from our paper and would form a completely new model, no longer simply adapting ControlNet as-is for our datasets.